# Enhancing Test Time Adaptation with Few-shot Guidance

### v1

@@ Introduction
$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (3)
@@
Fine-Tuning Source Model.
To enhance the adaptation of the pre-trained source model to the target, we employ the few-shot support set to fine-tune the model with the FDA module. Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (4)
@@
Table 2: Compared with test time adaptation methods on three datasets with ResNet50 backbone.
@@
ERM (Vapnik 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN (Nado et al. 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT (Wang et al. 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A (Iwasawa and Matsuo 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA (Niu et al. 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME (Boudiaf et al. 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
TSD (Wang, Zhang et al. 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
@@
FT+TENT (Wang et al. 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD (Wang, Zhang et al. 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
@@
Stage II: Test Time Adaptation
During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives. The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon 2001). Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Fig. 3.
Entropy Filter.
To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon 2001) to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as:
$H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (5)
Based on the insights from previous work (Wang et al. 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},...,\hat{x}_{|\alpha\%*B|}\}$.
Prototype Memory Bank.
We maintain a prototype memory bank $M=\{m_{1},m_{2},...,m_{C}\}$ to store class prototypes, where $C$ represents categories of the target. The prototype memory bank is initialized with the few-shot support set $S$, defined as:
$m_{c_{0}}=\frac{\sum_{i=1}^{|S|}f_{i}\cdot\mathbbm{1}[y_{i}=c]}{\sum_{i=1}^{|S|}\mathbbm{1}[y_{i}=c]},$ (6)
where $\mathbbm{1}[\cdot]$represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype. Thanks to the few-shot support, precise guidance can be provided during the initial phase, thereby reducing reliance on the quality of online mini-batch data.
$m_{c_{t}}=\beta*m_{c_{t-1}}+(1-\beta)*\frac{\sum_{j=1}^{|\hat{x}|}f_{j}\cdot\mathbbm{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{|\hat{x}|}\mathbbm{1}[\hat{y}_{j}=c]},$ (7)
Test Time Adaptation.
During the test time adaptation, we adopt high-quality pseudo-labeled samples to guide the model update. First, we define the prototype-based classification output as the softmax over the feature similarity to prototypes for class $c$:
@@
$\mathcal{M}_{j}=\mathbbm{1}\left[\arg\max p_{j}=\arg\max\hat{p}_{j}\right].$ (9)
@@
$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{|\hat{x}|}}\mathcal{M}_{j}}.$ (10)
@@
Experiment
Experimental Settings
Dataset.
To test the effectiveness of our setting and method, we experiment on three cross-domain benchmarks. PACS (Li et al. 2017) consists of four distinct domains: Art, Cartoon, Photo, and Sketch, and comprises a total of 9,991 images and 7 classes. Office-Home (Venkateswara et al. 2017) consists of 15,588 images from four distinct image domains, namely Real World, Clipart, Art, and Product, with 65 classes. DomainNet (Peng et al. 2019) is a large-scale dataset that comprises six domains, denoted as $d\in${Clipart, Infograph, Painting, Quickdraw, Real, Sketch}, with a total of 586,575 images and 345 classes.
Implementation Details.
In the main experiments, we utilize ResNet-50 (He et al. 2016) pre-trained on ImageNet-1k (Russakovsky et al. 2015) as our backbone, which is widely used in test time adaptation literature. For source model training, we adhere to the leave-one-domain-out protocol, as recommended by previous work  (Wang, Zhang et al. 2023; Zhou et al. 2021). We employ the Adam optimizer with a learning rate of $5e^{-5}$. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
Baselines.
We compare our method with test time adaptation methods, BN (Nado et al. 2020), TENT (Wang et al. 2021), ETA (Niu et al. 2022), T3A (Iwasawa and Matsuo 2021), LAME (Boudiaf et al. 2022), and TSD (Wang, Zhang et al. 2023). Additionally, we create new baselines by combining fine-tuning with TTA methods for a more comprehensive comparison. We also compare with some methods in domain generalization and source-free domain adaptation, including DNA (Chu et al. 2022), PCL (Yao et al. 2022), SWAD (Cha et al. 2021), and F-mix (Kundu et al. 2022).
@@
Performance Comparisons
Comparison with TTA methods.
Tab. 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Tab. 3). We observe that our method achieves state-of-the-art performance.
@@
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines, which combine fine-tuning with TTA methods (e.g., TENT and TSD). According to the results in Tab. 2, our method shows an average (three datasets) improvement of 4.2% over Fine-Tuning+TENT, and 2.4% over Fine-Tuning+TSD, which demonstrates the superiority of our method over existing TTA techniques in migrating to few-shot TTA scenarios. This result benefits from our unique FDA module design in Fine-Tuning and the use of few-shot support set to initialize prototype memory bank.
Refer to captionRefer to caption(a) Different stage(b) Different $\alpha$
@@
Figure 4: (a) effectiveness analysis about two-stage framework and (b) sensitivity analysis about parameter $\alpha$.
Comparison with DG/SFDA methods.
@@
The above experiments mainly focus on TTA, which aims to adapt the model during the test time. A natural question arises: How about our method compared with domain generalization (DG) or source-free domain adaptation (SFDA) methods?
@@
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Tab. 3. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy. This result benefits from our adaptation of the model during the test time. In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Tab. 3. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Ablation Study
Figure 5: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) our method.
Effectiveness of two-stage framework.
@@
Our proposed method consists of two stages, with the individual contributions of each stage presented in Fig. 4(a). Compared to the baseline source model, Stage I of our approach achieves an average improvement of 6.6% on the Office-Home. This highlights the effectiveness of our fine-tuning strategy, which employs a mixture of statistics between samples, validating its suitability for the target domain. Our test time adaptation method, which relies on class prototype memory bank guidance during Stage II, adds an extra 2.5% performance enhancement. As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
Sensitivity to $\alpha$.
@@
The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Fig. 4(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
Qualitative analysis by t-SNE visualization.
We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Fig. 5. The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Fig. 5(a). Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Fig. 5(b) and Fig. 5(c). In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Fig. 5(d).
Efficiency analysis.
In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Fig. 6(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Fig. 6(b), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
Refer to captionRefer to caption(a) Accuracy (batch size)(b) Running Time
Figure 6: During the test time adaptation, we analyze:(a) accuracy and (b) running time, across various batchsize.
Figure 7: Ablation experiments on shot size.
Ablation experiments on shot size.
@@
To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 7, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Conclusion

### v2

@@ 4.2 Stage I: Model Adaptation via Fine-Tuning
$f_{i}^{\prime}=\bm{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\bm{\beta}_{\operatorname{mix}},$ (7)
@@
Fine-Tuning Source Model. To enhance the adaptation of the pre-trained source model to the target, we employ the few-shot support set to fine-tune the model with the FDA module. Specifically, the few-shot support set is processed through $f_{\theta_{s}}$ to minimize a supervised loss, defined as:
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\bm{s}_{i}\right)\right),$ (8)
@@
4.3 Stage II: Test Time Adaptation
During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives. The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon, 2001). Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Figure 3.
Entropy Filter. To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon, 2001) to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as:
$H(p_{i})=-\sum(p_{i})\cdot log(p_{i}).$ (9)
Based on the insights from previous work (Wang et al, 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\alpha\cdot B\rfloor}\}$.
Prototype Memory Bank. We maintain a prototype memory bank $M=\{m_{1},m_{2},...,m_{C}\}$ to store class prototypes, where $C$ represents categories of the target. The prototype memory bank is initialized with the few-shot support set $S$, defined as:
$m_{c_{0}}=\frac{\sum_{i=1}^{\lvert S\rvert}f_{i}\cdot\mathds{1}[y_{i}=c]}{\sum_{i=1}^{\lvert S\rvert}\mathds{1}[y_{i}=c]},$ (10)
where $\mathds{1}[\cdot]$ represents an indicator function, yielding a value of 1 if the argument is true or 0 otherwise, and $m_{c_{0}}$ denotes the initial moment of the $c$-th class prototype. Thanks to the few-shot support, precise guidance can be provided during the initial phase, thereby reducing reliance on the quality of online mini-batch data.
$m_{c_{t}}=\beta\cdot m_{c_{t-1}}+(1-\beta)\cdot\frac{\sum_{j=1}^{\lvert\hat{x}\rvert}f_{j}\cdot\mathds{1}[\hat{y}_{j}=c]}{\sum_{j=1}^{\lvert\hat{x}\rvert}\mathds{1}[\hat{y}_{j}=c]}.$ (11)
Table 2: Table 2: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone. FS-TTA achieves consistent improvements over TSD (Wang et al, 2021), the strongest baseline method.
@@ 4.3 Stage II: Test Time Adaptation
ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN (Nado et al, 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT (Wang et al, 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A (Iwasawa and Matsuo, 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA (Niu et al, 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME (Boudiaf et al, 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
TSD (Wang et al, 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
PROGRAM (Sun et al, 2024) 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
DEYO (Lee et al, 2024) 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
@@
FT+TENT (Wang et al, 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD (Wang et al, 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
@@
Test Time Adaptation. During the test time adaptation, we adopt high-quality pseudo-labeled samples to guide the model update. First, we define the prototype-based classification output as the softmax over the feature similarity to prototypes for class $c$:
@@
$\mathcal{M}_{j}=\mathbbm{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
@@
$\mathcal{L}_{online}=\frac{{\textstyle\sum_{j=1}^{\|\hat{x}\|}}\mathcal{H}_{j}*\mathcal{M}_{j}}{{\textstyle\sum_{j=1}^{\|\hat{x}\|}}\mathcal{M}_{j}}.$ (14)
@@
5 Experiment
5.1 Experimental Settings
Dataset. To evaluate the effectiveness of our proposed setting and method, we conduct experiments on three cross-domain benchmarks.
  • PACS (Li et al, 2017) consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket). Each domain contains seven object categories: dog, elephant, giraffe, guitar, horse, house, and person.
  • Office-Home (Venkateswara et al, 2017) comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
  • DomainNet (Peng et al, 2019) is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
Implementation Details. In our main experiments, we employ ResNet-50(He et al, 2016), pre-trained on ImageNet-1k(Russakovsky et al, 2015), as the backbone model, as it is widely adopted in the test-time adaptation literature. For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies (Wang et al, 2023; Zhou et al, 2021), treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5. Both the dropout probability and weight decay are set to zero. The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in (Cha et al, 2021). All images are resized to 224 × 224, and data augmentation is applied during source domain training, including random cropping, horizontal flipping, color jittering, and intensity adjustments. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
Figure 4: Comprehensive comparison between our method and the state-of-the-art method in DG/TTA settings on DomainNet.
@@ 5.1 Experimental Settings
Baselines. We compare our method with various test-time adaptation (TTA) approaches, including BN(Nado et al, 2020), TENT(Wang et al, 2021), ETA(Niu et al, 2022), T3A(Iwasawa and Matsuo, 2021), LAME(Boudiaf et al, 2022), TSD(Wang et al, 2023),PROGRAM (Sun et al, 2024) and DEYO (Lee et al, 2024). Additionally, we establish new baselines by integrating fine-tuning with existing TTA methods to ensure a more comprehensive comparison. Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA(Chu et al, 2022), PCL(Yao et al, 2022), SWAD(Cha et al, 2021), and F-mix(Kundu et al, 2022). Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN (Li et al, 2016), $L^{2}$ (Li et al, 2018a), DELTA (Li et al, 2019), FLUTE (Triantafillou et al, 2021), LCCS (Zhang et al, 2022). For a global overview, we compare our method with state-of-the-art method in various settings, as shown in Figure 4.
5.2 Performance Comparisons
Comparison with TTA methods. Table 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table 3). We observe that our method achieves state-of-the-art performance.
@@ 5.2 Performance Comparisons
Table 4: Compared with few-shot transfer learning methods on PACS dataset.
PACS
Method Art Cart Phot Sket Avg.
Few-shot transfer learning methods
AdaBN (Li et al, 2016) 85.0 83.5 96.0 78.7 85.8
$L^{2}$ (Li et al, 2018a) 85.6 84.1 96.4 76.3 85.6
DELTA (Li et al, 2019) 85.6 83.8 96.5 76.3 85.6
FLUTE (Triantafillou et al, 2021) 87.2 86.1 97.2 81.7 88.1
LCCS (Zhang et al, 2022) 87.7 86.9 97.5 83.0 88.8
FS-TTA 90.4 89.8 97.6 87.9 91.4
@@
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines by combining fine-tuning with representative TTA approaches. Specifically, we select TENT (Wang et al, 2021), as a widely adopted and foundational method in test-time adaptation, and TSD (Wang et al, 2023), which demonstrates state-of-the-art performance across benchmarks. According to the results in Table 2, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets. These results highlight the superiority of our framework in migrating to the few-shot TTA setting, benefiting from the proposed FDA module and the support-set-based prototype initialization.
@@
Figure 5: Ablation Study on (a) effectiveness analysis about two-stage framework and (b) effectiveness analysis about FDA module.
Figure 6: Comprehensive Sensitivity Analysis of (1) Different Entropy Filter Proportion $\alpha$ and K-shot Number on Model Performance.
@@
Comparison with DG/SFDA methods. The above experiments mainly focus on TTA, which aims to adapt the model during the test time. A natural question arises: How about our method compared with domain generalization (DG) or source-free domain adaptation (SFDA) methods?
@@
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Table 3. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy. This result benefits from our adaptation of the model during the test time. In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Table 3. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Comparison with few-shot transfer learning methods. In our research, we focus on Few-Shot Test Time Adaptation (FS-TTA), which utilizes a small number of target domain samples to enhance adaptation. To ensure a comprehensive evaluation, we compare our approach with existing few-shot transfer learning methods. The results on the PACS dataset are presented in Table 4. According to the results, FS-TTA consistently outperforms all baseline methods across different domains, achieving the highest average accuracy of 91.4%, which surpasses the best-performing baseline LCCS (88.8%). This result highlights the effectiveness of our approach in adapting to domain shifts and improving classification performance in the few-shot setting.
5.3 Ablation Study
@@
Effectiveness of two-stage framework. Our proposed method consists of two stages, with the individual contributions of each stage presented in Figure 5(a). Compared to the baseline source model, Stage I of our approach achieves an average improvement of 6.6% on the Office-Home. This highlights the effectiveness of our fine-tuning strategy, which employs a mixture of statistics between samples, validating its suitability for the target domain. Our test time adaptation method, which relies on class prototype memory bank guidance during Stage II, adds an extra 2.5% performance enhancement. As a result, our two-stage framework establishes itself as a robust foundation for the Few-Shot Test Time Adaptation setting, demonstrating its considerable potential in enabling online model adaptation in real-world situations where labeled data is scarce.
Figure 7: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) FS-TTA (Ours).
Figure 8: Analyzing the Impact of Batch Size on Accuracy and Running Efficiency.
Effectiveness of FDA module. In our first phase, we introduce the FDA module to tackle overfitting issues through feature augmentation. Here we conduct additional ablation experiments on the FDA module and compare it with Mix-up augmentation, as depicted in Figure 5(b). The results from the ablation experiments indicate that the FDA module is effective and outperforms mix-up augmentation. The baseline method (without any techniques in the fine-tuning phase) achieves an accuracy of 76.12%, while incorporating Mix-up leads to a slight improvement, reaching 76.2% (+0.08%). However, when our FDA module is introduced, the performance further increases to 76.49%, yielding a notable improvement of +0.37% over the baseline. These results highlight the advantage of FDA module in enhancing feature diversity and robustness, surpassing standard augmentation techniques like Mix-up.
@@
Sensitivity to $\alpha$. The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Figure 6(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
@@
Ablation experiments on shot size. To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 6, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Qualitative analysis by t-SNE visualization. We present t-SNE visualizations to compare the feature representations of the pre-trained source model (ERM), test time adaptation methods (TENT and LAME), and our proposed method, as illustrated in Figure 7. The learned features of the pre-trained source model on the target domain are not well-separated due to the significant domain gap, as shown in Figure 7(a). Additionally, we can observe no considerable feature distribution changes on the target domain after adaptation with TENT and LAME methods, as shown in Figure 7(b) and Figure 7(c). In contrast, our method produces more uniform and aligned feature distribution after adapting to the target domain, as shown in Figure 7(d).
Efficiency analysis. In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Figure 8(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Figure 8(b), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
Table 5: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
Backbones PACS Office-Home
ResNet 84.59 67.37
+ TSD (Wang et al, 2023) 89.41 68.67
+ Ours 91.42 76.49
ViT-B/16 87.13 79.06
+ TSD (Wang et al, 2023) 90.20 81.80
+ Ours 91.89 87.32
Scalability on Vision Transformer. We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) (Dosovitskiy et al, 2021). Specifically, we adopt ViT-B/16 as the backbone and compare the baseline TSD with our approach. The results, shown in Table 5, demonstrate that our method achieves consistent improvements over TSD. On the PACS dataset, our method improves the accuracy from 90.20% (TSD) to 91.89%, while on Office-Home, it further boosts performance from 81.80% to 87.32%. These gains highlight that our approach is not limited to convolutional networks but can also optimize transformer-based architectures, making it a versatile solution for various backbone choices.
6 Conclusion